four times the measure of an angle is 15 less than the measure of its compliment....
I tried the equation: <2*4+15=<1
Please help me!!! thank you!!

Answer by checkley77(12844)

4X=(90-X)-15
4X=75-X
4X+X=75
5X=75
X=75/5
X=15 DEGREES IS THE SMALLER ANGLE.
90-15=75 DEGREES FOR THE COMPLIMENTARY ANGLE.
PROOF:
4*15=75-15
60=60